# CHARM: Considering Human Attributes for Reinforcement Modeling

### v1

@@ I Introduction
Reinforcement Learning from Human Feedback (RLHF) offers a solution for robots to learn and adapt their policies to human preferences. Despite much previous work in learning or simulating human feedback has mentioned that human feedback patterns differ from person to person and may be affected by factors beyond the task itself [1, 2, 3, 4], most prior work assumes that all human feedback is correct, or directly simulates human feedback from reward values with random noises [5]. RLHF algorithms require a substantial amount of high-quality human feedback for learning. Oversimplifying human feedback patterns will introduce a lab-to-real-world discrepancy and result in a performance drop for the learning algorithms. Our key insight is that human feedback is not only associated with the task and the agent’s performance, but also affected by human characteristics. In this work, we therefore concern ourselves with the problem of investigating the relation between human feedback patterns and human characteristics.
@@
RLHF has gained significant attention in recent advancements in robot learning and large model post-training. However, RLHF methods are sensitive to data quality and changes in human feedback patterns [6]. It is widely understood that people’s reactions and responses to various situations are influenced by many factors, such as personality [7] and rationality level [4]. Human factors like teaching style, teaching experience, educational background, robot experience, and trust in robots, though not definitively linked, could play a role in shaping how individuals provide feedback in the context of teaching robots [3]. Real human feedback is often biased [8], erroneous [9], delayed [10], and sometimes even missed [11]. Different individuals may also give varying feedback for the same observation [12]. Furthermore, RLHF often requires a large amount of human feedback, which can be costly to collect. One solution for mitigating the data shortage and reducing the cost is to use simulated human feedback generated by oracles. Commonly used oracles, often based on optimal policies [13][14] or heuristic functions [15][16][17], did not account for human factors.
Despite the recognized importance and broad existence of these dynamic human feedback patterns, few studies have explicitly considered how factors beyond the task affect human feedback [2]. Assuming users would consistently provide feedback only based on the task or relying on oracles to provide instant, bias-free feedback oversimplifies how humans interact with robots [5]. These oversimplifications can result in biased outcomes when transitioning from in-lab developments to real world applications. To bridge this gap, our work focuses on investigating the impact of human characteristics on feedback patterns and exploring if incorporating these attributes can improve predicting human feedback.
@@
In this work, we demonstrated that our selected human characteristics exert a limited correlation with feedback delay but a pronounced correlation with feedback values. While it is well-established that human feedback patterns can vary among individuals [1, 18], to the best of our knowledge, this is the first work that systematically investigated how human characteristics are associated with human feedback and leveraged human characteristics to predict human teaching signals. We believe that this work has the potential to reduce the gap between robot learning in the lab and robot learning in the wild and benefit the development of RLHF methods.
@@ II Background
Reinforcement Learning from Human Feedback. Reinforcement Learning from Human Feedback (RLHF) is a technique that enables agents to learn and adapt their behavior based on feedback provided by humans, rather than relying solely on predefined reward functions [19]. RLHF has been applied successfully in various fields, particularly in robot learning, where it allows robots to adjust their strategies based on human preferences [20] [15], critiques [21], and ratings [22], and natural language [23]. Research indicates that human feedback can significantly reduce unnecessary exploration during the early training stages [24], leading to more efficient convergence [10].
@@
Human Oracles in RLHF. One major challenge of RLHF is the large volume of human feedback required [25]. Using oracles can significantly reduce the need for human feedback, lowering costs associated with participant recruitment and briefings[26] [27]. One common approach is to generate oracles from heuristic functions. In this method, a hand-engineered function maps state-action pairs to positive or negative critiques instead of relying on binary feedback directly from users [28] [29] [30]. Another approach generates oracles from optimal policies, where an optimal policy is usually a fully trained model and feedback is given based on the alignment between the current robot policy and the optimal policy [17] [21]. However, despite their convenience, oracles can deviate from real human feedback [31]. Human feedback can be biased [32], unreliable[33], and inconsistent [34, 1]. Furthermore, human feedback is rarely instantaneous; delays can occur due to distraction [35], or hesitation [36]. Relying on perfect oracles that provide instant, error-free feedback oversimplifies the complexity of human-robot interactions [5]. Moreover, human feedback patterns can vary across individuals and often cannot be captured by perfect oracles, which creates a gap between in-lab performance and real-world performance.
@@
Human Characteristics and Human Feedback Patterns. Personality plays a role in shaping how people interact with and respond to different situations [7][37]. The widely used Big Five personality model—comprising openness, conscientiousness, extraversion, agreeableness, and neuroticism—captures these traits [38]. Individuals with higher conscientiousness are efficient and responsible [39] and may provide less delayed feedback. People high in neuroticism may deliver high-quality feedback [40], especially during intensive robotic tasks. In addition to personality, teaching experience and teaching style may also indirectly affect human teaching behavior. People with formal teaching experience may provide more structured and constructive feedback [41] [42]. Educational background, robot experience, and trust in robots further shape feedback patterns. Individuals with more robot experience or with relevant technical backgrounds may better understand the nuances of interaction with robots [43], providing more precise feedback. Trust in robots could also play a role—those with a higher trust may be more forgiving of robot errors [44], resulting in biased feedback.
@@ III Methodology
Providing feedback is a complex process involving multiple cognitive functions [45]. We argue that human feedback does not solely rely on tasks or robot behaviors, it also depends on the human teachers themselves. To explore this, we developed a questionnaire to capture six key domains of human characteristics and conducted a public study to collect both human feedback and their characteristics. Moreover, we proposed a novel method for training human feedback simulator CHARM, Considering Human Attributes for Reinforcement Modeling.
@@ III-A Human Feedback pattern
Real human feedback differs from simulated human feedback in two key aspects: subjective evaluation values and the timing of feedback delivery. Scalar feedback offers a rich representation by capturing both positive and negative magnitudes while incorporating delay accounts for individual differences rather than assuming a uniform delay distribution [10]. Thus, in this work, we represent human feedback patterns using scalar feedback and delays.
@@ III-B Questionnaire
As mentioned in prior work [7, 37, 38, 39, 40], human feedback patterns can vary from person to person. In this work, we designed a questionnaire to capture the differences in human characteristics among individuals. The questionnaire consists of six domains, including trust in robots, robot experience, educational background, teaching experience, teaching style, and personality traits. The details of the questionnaire we used are as follows:
  • Trust in Robots: We combined two questions from [46] and one from [47] to measure participants’ trust in robots, focusing on perceived reliability, safety, and concerns about overreliance.
  • Robot Experience: We adapted two questions from [48] to evaluate participants’ familiarity with robots, emphasizing their prior experiences with both general robots and robotic arms.
@@
  • Personality Traits: We employed a 10-item short version of the Big Five Inventory [49] to assess individual personality differences, covering openness, conscientiousness, extraversion, agreeableness, and neuroticism.
  • Teaching Style: We adapted eight questions from [50] to capture participants’ teaching styles in classroom scenarios—drawing on their actual teaching experiences when available or assuming their approach otherwise—with a focus on relationship building, routine setting, care, praise, and behavior monitoring.
@@ III-C Considering Human Attributes for Reinforcement Modeling
1 Pre-train agents for feedback collection
@@ IV-A Task Setup
Our experiment used two tasks from mimicgen [51], nut assembly, and coffee preparation, as shown in Figure 3. For each task, we pre-trained agents with different steps using the BC-RNN algorithm [52], leveraging 1,000 human expert demonstrations from [51]. For each task, we trained three agents, 500 epochs (minimally trained), 1,000 epochs (medium-trained), and 2,000 epochs (well-trained), to generate trajectories with diverse proficiencies. During the experiment, participants were shown six randomly sampled trajectories (one trajectory per agent, three agents for two tasks). Additionally, we used the Density-Based Reward Modeling algorithm [53] to extract the reward function from expert demonstrations, which provided rewards and served as the task statistics for each task.
@@ IV-B GUI for Standardized Feedback Collection
We developed a graphical user interface (GUI) to standardize and streamline data collection. The GUI consists of two components: the ”state window” and the ”noticing window”. During robot execution, the state window remained continuously displayed. Upon completing a task, the noticing window appeared with a blinking prompt, instructing the participant to provide a rating. This window displayed five emoji buttons labeled from ”Very Dissatisfied” to ”Very Satisfied,” corresponding to a five-point scale in a range of $[-2,2]$. Additionally, the system recorded the response latency as the elapsed time between the onset of the noticing window and the participant’s rating submission, thereby capturing feedback delay. Once a rating was submitted, the noticing window disappeared, and the state window was reactivated for the subsequent trajectory execution. If no response was provided, the noticing window remained visible for a maximum of 5 seconds before automatically proceeding to the next trajectory.
@@ IV-C Procedure
Figure 3: Robotic tasks from mimicgen [51] in robosuite simulator[54].
@@ V-A Collected Human Feedback and Questionnaire Responses
As shown in the top row, participants’ delays are tightly clustered around one second, which supports the conclusion in prior work [55]. The reward distribution broadly aligns with the human feedback distribution. We computed feedback accuracy for each participant by considering $\{-2,-1\}$ as negative, $\{1,2\}$ as positive, and $\{0\}$ as neutral. On average, participants achieved an accuracy of approximately $51.79\%$ with a standard deviation of $10.90\%$, and the highest individual accuracy reached $74.50\%$. We computed the Pearson correlation coefficient [56] between the reward and human feedback values and delays, respectively, yielding
$\rho_{r,value}\approx 0.563\quad\text{and}\quad\rho_{r,delay}\approx 0.0004,$
@@ V-C Including Human Characteristics Improves Feedback Prediction Accuracy
CHARM Training Setup As noted in Section V-A, our dataset shows an unbalanced distribution in feedback values. To improve the training quality, we applied a Box-Cox transformation [57] to the delay data and applied a weighted loss function during training to assign higher weights to minority classes in the feedback value distribution. We performed k-fold cross-validation (in this work, k=10) by randomly splitting our $4655$ data points into 90% training and 10% validation sets, with a new random seed every time. We demonstrated the models’ performance on two feedback value scales. One is the original value scale, a five-point scale, ranging from -2 to 2. The other one is a widely used feedback value scale, binary feedback, with only positive and negative. We adapted the data set for training the models to predict binary feedback by classifying $-2$ and $-1$ as negative, classifying $1$ and $2$ as positive, and dropping the 0s. We had three methods for predicting feedback values. Random prediction: the model randomly selected a feedback value (expected accuracy is 20% and 50%). Task statistics only: the model was trained to predict feedback values using corresponding task statistics. Human characteristics and task statistics (CHARM): the model was trained to predict feedback values using both corresponding task statistics and human characteristics.
Figure 6: Comparison of feedback prediction accuracy. Left: Feedback prediction on a five-point scale (ranging from -2 to 2). Right: Binary feedback prediction (positive vs. negative). Each bar represents the average accuracy across a 10-fold cross-validation. Results indicate that models trained with human characteristics and task statistics combined (CHARM) significantly outperform models trained with task statistics only ($p<0.001$).

### v2

@@ I Introduction
Reinforcement Learning from Human Feedback (RLHF) offers a solution for robots to learn and adapt their policies to human preferences. Despite much previous work in learning or simulating human feedback has mentioned that human feedback patterns differ from person to person and may be affected by factors beyond the task itself [yu2023thumbs, huang2024modeling, PANADERO2022100416, ghosal2023effectmodelinghumanrationality], most prior work assumes that all human feedback is correct, or directly simulates human feedback from reward values with random noises [krening2018interaction]. RLHF algorithms require a substantial amount of high-quality human feedback for learning. Oversimplifying human feedback patterns will introduce a lab-to-real-world discrepancy and result in a performance drop for the learning algorithms. Our key insight is that human feedback is not only associated with the task and the agent’s performance, but also affected by human characteristics. In this work, we therefore concern ourselves with the problem of investigating the relation between human feedback patterns and human characteristics.
@@
RLHF has gained significant attention in recent advancements in robot learning and large model post-training. However, RLHF methods are sensitive to data quality and changes in human feedback patterns [10161105]. It is widely understood that people’s reactions and responses to various situations are influenced by many factors, such as personality [bolmont2001relationships] and rationality level [ghosal2023effectmodelinghumanrationality]. Human factors like teaching style, teaching experience, educational background, robot experience, and trust in robots, though not definitively linked, could play a role in shaping how individuals provide feedback in the context of teaching robots [PANADERO2022100416]. Real human feedback is often biased [ibarz2018reward], erroneous [brawer2023interactive], delayed [knox2009interactively], and sometimes even missed [dromnelle2020coping]. Different individuals may also give varying feedback for the same observation [kupcsik2018learning]. Furthermore, RLHF often requires a large amount of human feedback, which can be costly to collect. One solution for mitigating the data shortage and reducing the cost is to use simulated human feedback generated by oracles. Commonly used oracles, often based on optimal policies [sahni2016policy][swamy2024minimaximalist] or heuristic functions [griffith2013policy][du2019provably][sheidlower2022keeping], did not account for human factors.
Despite the recognized importance and broad existence of these dynamic human feedback patterns, few studies have explicitly considered how factors beyond the task affect human feedback [huang2024modeling]. Assuming users would consistently provide feedback only based on the task or relying on oracles to provide instant, bias-free feedback oversimplifies how humans interact with robots [krening2018interaction]. These oversimplifications can result in biased outcomes when transitioning from in-lab developments to real world applications. To bridge this gap, our work focuses on investigating the impact of human characteristics on feedback patterns and exploring if incorporating these attributes can improve predicting human feedback.
@@
In this work, we demonstrated that our selected human characteristics exert a limited correlation with feedback delay but a pronounced correlation with feedback values. While it is well-established that human feedback patterns can vary among individuals [yu2023thumbs, christiano2017deep], to the best of our knowledge, this is the first work that systematically investigated how human characteristics are associated with human feedback and leveraged human characteristics to predict human teaching signals. We believe that this work has the potential to reduce the gap between robot learning in the lab and robot learning in the wild and benefit the development of RLHF methods.
@@ II Background
Reinforcement Learning from Human Feedback. Reinforcement Learning from Human Feedback (RLHF) is a technique that enables agents to learn and adapt their behavior based on feedback provided by humans, rather than relying solely on predefined reward functions [sutton2018reinforcement, yu2024much]. RLHF has been applied successfully in various fields, particularly in robot learning, where it allows robots to adjust their strategies based on human preferences [liu2022task] [griffith2013policy] [TAN2020113043], critiques [cederborg2015policy], and ratings [arumugam2019deep], and natural language [kuhlmann2004guiding]. Research indicates that human feedback can significantly reduce unnecessary exploration during the early training stages [cruz2015interactive], leading to more efficient convergence [knox2009interactively].
@@
Human Oracles in RLHF. One major challenge of RLHF is the large volume of human feedback required [corrado2024guided, yu2021active]. Using oracles can significantly reduce the need for human feedback, lowering costs associated with participant recruitment and briefings[wang2020teaching] [reddy2018shared]. One common approach is to generate oracles from heuristic functions. In this method, a hand-engineered function maps state-action pairs to positive or negative critiques instead of relying on binary feedback directly from users [martins2013heuristically] [rosenfeld2018leveraging] [mehta2024unified]. Another approach generates oracles from optimal policies, where an optimal policy is usually a fully trained model and feedback is given based on the alignment between the current robot policy and the optimal policy [sheidlower2022keeping] [cederborg2015policy]. However, despite their convenience, oracles can deviate from real human feedback [lindner2022humans]. Human feedback can be biased [thomaz2006reinforcement], unreliable[faulkner2020interactive], and inconsistent [bignold2021evaluation, yu2023thumbs]. Furthermore, human feedback is rarely instantaneous; delays can occur due to distraction [kessler2019active], or hesitation [loftin2014strategy]. Relying on perfect oracles that provide instant, error-free feedback oversimplifies the complexity of human-robot interactions [krening2018interaction]. Moreover, human feedback patterns can vary across individuals and often cannot be captured by perfect oracles, which creates a gap between in-lab performance and real-world performance.
@@
Human Characteristics and Human Feedback Patterns. Personality plays a role in shaping how people interact with and respond to different situations [bolmont2001relationships][1513803]. The widely used Big Five personality model—comprising openness, conscientiousness, extraversion, agreeableness, and neuroticism—captures these traits [soldz1999big]. Individuals with higher conscientiousness are efficient and responsible [mccrae1992introduction] and may provide less delayed feedback. People high in neuroticism may deliver high-quality feedback [pelgrim2014factors], especially during intensive robotic tasks. In addition to personality, teaching experience and teaching style may also indirectly affect human teaching behavior. People with formal teaching experience may provide more structured and constructive feedback [georgountzou2019peer] [brinko1993practice]. Educational background, robot experience, and trust in robots further shape feedback patterns. Individuals with more robot experience or with relevant technical backgrounds may better understand the nuances of interaction with robots [8525515], providing more precise feedback. Trust in robots could also play a role—those with a higher trust may be more forgiving of robot errors [karli2023if], resulting in biased feedback.
@@ III Methodology
Providing feedback is a complex process involving multiple cognitive functions [maior2018workload]. We argue that human feedback does not solely rely on tasks or robot behaviors, it also depends on the human teachers themselves. To explore this, we developed a questionnaire to capture six key domains of human characteristics and conducted a public study to collect both human feedback and their characteristics. Moreover, we proposed a novel method for training human feedback simulator CHARM, Considering Human Attributes for Reinforcement Modeling.
@@ III-A Human Feedback pattern
Real human feedback differs from simulated human feedback in two key aspects: subjective evaluation values and the timing of feedback delivery. Scalar feedback offers a rich representation by capturing both positive and negative magnitudes while incorporating delay accounts for individual differences rather than assuming a uniform delay distribution [knox2009interactively]. Thus, in this work, we represent human feedback patterns using scalar feedback and delays.
@@ III-B Questionnaire
As mentioned in prior work [bolmont2001relationships, 1513803, soldz1999big, mccrae1992introduction, pelgrim2014factors], human feedback patterns can vary from person to person. In this work, we designed a questionnaire to capture the differences in human characteristics among individuals. The questionnaire consists of six domains, including trust in robots, robot experience, educational background, teaching experience, teaching style, and personality traits. The details of the questionnaire we used are as follows:
  • Trust in Robots: We combined two questions from [charalambous2016development] and one from [nomura2004psychology] to measure participants’ trust in robots, focusing on perceived reliability, safety, and concerns about overreliance.
  • Robot Experience: We adapted two questions from [schaefer2013trust] to evaluate participants’ familiarity with robots, emphasizing their prior experiences with both general robots and robotic arms.
@@
  • Personality Traits: We employed a 10-item short version of the Big Five Inventory [rammstedt2007measuring] to assess individual personality differences, covering openness, conscientiousness, extraversion, agreeableness, and neuroticism.
  • Teaching Style: We adapted eight questions from [ertesvaag2011measuring] to capture participants’ teaching styles in classroom scenarios—drawing on their actual teaching experiences when available or assuming their approach otherwise—with a focus on relationship building, routine setting, care, praise, and behavior monitoring.
@@ III-C Considering Human Attributes for Reinforcement Modeling
1Pre-train agents for feedback collection
@@ IV-A Task Setup
Our experiment used two tasks from mimicgen [mandlekar2023mimicgen], nut assembly, and coffee preparation, as shown in Figure 3. For each task, we pre-trained agents with different steps using the BC-RNN algorithm [mandlekar2021matters], leveraging 1,000 human expert demonstrations from [mandlekar2023mimicgen]. For each task, we trained three agents, 500 epochs (minimally trained), 1,000 epochs (medium-trained), and 2,000 epochs (well-trained), to generate trajectories with diverse proficiencies. During the experiment, participants were shown six randomly sampled trajectories (one trajectory per agent, three agents for two tasks). Additionally, we used the Density-Based Reward Modeling algorithm [gleave2022imitation] to extract the reward function from expert demonstrations, which provided rewards and served as the task statistics for each task.
@@ IV-B GUI for Standardized Feedback Collection
We developed a graphical user interface (GUI) to standardize and streamline data collection. The GUI consists of two components: the ”state window” and the ”noticing window”. During robot execution, the state window remained continuously displayed. Upon completing a task, the noticing window appeared with a blinking prompt, instructing the participant to provide a rating. This window displayed five emoji buttons labeled from ”Very Dissatisfied” to ”Very Satisfied,” corresponding to a five-point scale in a range of [2,2][-2,2]. Additionally, the system recorded the response latency as the elapsed time between the onset of the noticing window and the participant’s rating submission, thereby capturing feedback delay. Once a rating was submitted, the noticing window disappeared, and the state window was reactivated for the subsequent trajectory execution. If no response was provided, the noticing window remained visible for a maximum of 5 seconds before automatically proceeding to the next trajectory.
@@ IV-C Procedure
Figure 3: Robotic tasks from mimicgen [mandlekar2023mimicgen] in robosuite simulator[zhu2020robosuite].
@@ V-A Collected Human Feedback and Questionnaire Responses
As shown in the top row, participants’ delays are tightly clustered around one second, which supports the conclusion in prior work [hockley1984analysis]. The reward distribution broadly aligns with the human feedback distribution. We computed feedback accuracy for each participant by considering $\{-2,-1\}$ as negative, $\{1,2\}$ as positive, and $\{0\}$ as neutral. On average, participants achieved an accuracy of approximately $51.79\%$ with a standard deviation of $10.90\%$, and the highest individual accuracy reached $74.50\%$. We computed the Pearson correlation coefficient [schober2018correlation] between the reward and human feedback values and delays, respectively, yielding
$\rho_{r,\text{value}}\approx 0.563,p<0.001;\quad\rho_{r,\text{delay}}\approx-0.004,p=0.769.$
@@ V-C Including Human Characteristics Improves Feedback Prediction Accuracy
CHARM Training Setup As noted in Section V-A, our dataset shows an unbalanced distribution in feedback values. To improve the training quality, we applied a Box-Cox transformation [SIBONO2025343704] to the delay data and applied a weighted loss function during training to assign higher weights to minority classes in the feedback value distribution. We performed k-fold cross-validation (in this work, k=10) by randomly splitting our $4655$ data points into 90% training and 10% validation sets, with a new random seed every time. We demonstrated the models’ performance on two feedback value scales. One is the original value scale, a five-point scale, ranging from -2 to 2. The other one is a widely used feedback value scale, binary feedback, with only positive and negative. We adapted the data set for training the models to predict binary feedback by classifying 2-2 and 1-1 as negative, classifying $1$ and $2$ as positive, and dropping the 0s. We had three methods for predicting feedback values. Random prediction: the model randomly selected a feedback value (expected accuracy is 20% and 50%). Task statistics only: the model was trained to predict feedback values using corresponding task statistics. Human characteristics and task statistics (CHARM): the model was trained to predict feedback values using both corresponding task statistics and human characteristics.
Figure 6: Comparison of feedback prediction accuracy. Left: Feedback prediction on a five-point scale (ranging from -2 to 2). Right: Binary feedback prediction (positive vs. negative). Each bar represents the average accuracy across a 10-fold cross-validation. Results indicate that models trained with human characteristics and task statistics combined (CHARM) significantly outperform models trained with task statistics only ($p<0.001$, for all paired $t$-tests).